INFRARED DIODE LASER ABSORPTION SPECTROSCOPY OF ^{1}A′ HC^{35}Cl

BOR-CHEN CHANG, Department of Chemistry, National Central University, Chung-Li, Taiwan 32054; \underline{TREVOR J. SEARS} and RUIAN FEI, Department of Chemistry, Brookhaven National Laboratory, Upton, NY, 11973-5000.

The C-Cl stretching fundamental of HC^{35}Cl in its ground, ^{1}A′, state has been detected by transient diode laser absorption spectroscopy. The radical was formed by the excimer laser photolysis of CHBr_{2}Cl at 193 nm in a continuous flow system. The band has a-type character and the fundamental vibrational frequency is 811.596 cm^{−1} . No evidence was found for perturbations in either the ground or vibrationally excited levels, implying that the excited ã ^{3}A′′ state lies above the CCl stretching level in HCCl. The observed vibration-rotation transitions were well fit by a standard asymmetric rotor Hamiltonian model. There is a decrease in the A rotational constant on vibrational excitation, suggesting some bend-stretch coupling is present.

Acknowledgments: This work was carried out at Brookhaven National Laboratory under Contract No. DE-AC02-76CH00016 with the U. S. Department of Energy and supported by its Division of Chemical Sciences, Office of Basic Energy Sciences. BCC also thanks the National Science Council, Taiwan for its support of this work under Grant NSC86-2112-M-008-032-T.